# GHPO: Adaptive Guidance for Stable and Efficient LLM Reinforcement Learning

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has recently emerged as a powerful paradigm for facilitating the self-improvement of large language models (LLMs), particularly in the domain of complex reasoning tasks. However, prevailing on-policy RL methods often contend with significant training instability and inefficiency. This is primarily due to a capacity-difficulty mismatch, where the complexity of training data frequently outpaces the model’s current capabilities, leading to critically sparse reward signals and stalled learning progress. This challenge is particularly acute for smaller, more resource-efficient LLMs. To overcome this, we introduce the Guided Hybrid Policy Optimization (GHPO), a novel difficulty-aware reinforcement learning framework. GHPO dynamically calibrates task difficulty by employing adaptive prompt refinement to provide targeted guidance. This unique approach adaptively balances direct imitation learning for problems currently beyond the model’s reach with exploration-based reinforcement learning for more manageable tasks, effectively creating a smooth and optimized learning curriculum. Extensive experiments demonstrate that GHPO achieves an average performance gain of approximately 5% across six challenging mathematics benchmarks, consistently outperforming strong on-policy reinforcement learning and curriculum learning baselines. Further analysis confirms that our framework significantly enhances both training stability and final reasoning performance, thus offering a scalable and efficient solution for developing powerful and robust reasoning models.
Figure 1: Overall Performance of GHPO across six benchmarks on the Qwen2.5-7B base model.
@@ 1 Introduction
In this work, we introduce a simple yet effective solution: guiding the model with partial ground truth solution traces. By conditioning the model on these traces, we steer its output distribution closer to the correct answer, which alleviates the reward sparsity problem for difficult samples. However, naively applying this technique risks making the training data too easy, reducing learning efficiency on problems the model could have solved on its own. To address this, we propose the Guided Hybrid Policy Optimization (GHPO) framework. GHPO incorporates a dynamic mechanism that first assesses sample difficulty and then adaptively applies different level of guidance. For problems deemed within the model’s reach, it employs standard on-policy RL to encourage exploration. For overly challenging samples, it seamlessly transitions to a form of imitation learning by providing solution traces. This hybrid approach automatically balances exploration and imitation learning, preserving training efficiency for manageable tasks while effectively providing adaptive guidance for difficult ones, thereby enhancing both training stability and sample efficiency.
@@
  • We identify the critical role of capacity alignment in RLVR and propose the GHPO framework to improve training stability and efficiency.
  • We introduce an automated framework to detect sample difficulty and adaptively switch between on-policy reinforcement learning and guided imitation learning by adaptive prompt refinement.
@@ 3.1 From Static Guidance to a Dynamic Framework
As established, our core strategy is to integrate guidance directly into the reinforcement learning loop, conditioning the policy on partial ground-truth traces to overcome the reward sparsity detailed in Section 2.3. This approach is motivated by Assumption 1, which posits that such guidance increases the likelihood of success on difficult problems, thereby providing a valid learning signal where one would otherwise be absent.
@@ 4.3 Implementation Details
We conducted our experiments using the openr1 [17] codebase and the TRL [18] framework. GHPO training was performed on 8 powerful GPUs, each equipped with 80GB memory and high memory bandwidth. The implementation code is available online to ease reproducibility [^1] https://github.com/hkgc-1/GHPO.

### v2

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has recently emerged as a powerful paradigm for facilitating the self-improvement of large language models (LLMs), particularly in the domain of complex reasoning tasks. However, prevailing on-policy RL methods often contend with significant training instability and inefficiency. This is primarily due to a capacity-difficulty mismatch, where the complexity of training data frequently outpaces the model’s current capabilities, leading to critically sparse reward signals and stalled learning progress. This challenge is particularly acute for smaller, more resource-efficient LLMs. To overcome this, we introduce the Guided Hybrid Policy Optimization (GHPO), a novel difficulty-aware reinforcement learning framework. GHPO dynamically calibrates task difficulty by employing adaptive prompt refinement to provide targeted guidance. This unique approach adaptively balances direct imitation learning for problems currently beyond the model’s reach with exploration-based reinforcement learning for more manageable tasks, effectively creating a smooth and optimized learning curriculum. Extensive experiments demonstrate that GHPO achieves an average performance gain of approximately 5% across six challenging mathematics benchmarks, consistently outperforming strong on-policy reinforcement learning and curriculum learning baselines. Further analysis confirms that our framework significantly enhances both training stability and final reasoning performance, thus offering a scalable and efficient solution for developing powerful and robust reasoning models. The implementation code is available online to ease reproducibility [^1] https://github.com/hkgc-1/GHPO.
Figure 1: Overall Performance of GHPO across several benchmarks on the Qwen2.5-7B base model.
@@ 1 Introduction
In this work, drawing inspiration from imitation learning techniques like SFT, we introduce a simple yet effective solution: guiding the model with partial ground truth solution traces. By conditioning the model on these traces, we steer its output distribution closer to the correct answer, which alleviates the reward sparsity problem for difficult samples. However, a naive application of this technique risks making the training data too easy, potentially reducing learning efficiency on problems the model could have solved independently. To address this, we propose the Guided Hybrid Policy Optimization (GHPO) framework. GHPO ingeniously combines online Reinforcement Learning (RL) and imitation learning within a unified framework. It uses a dynamic mechanism to first assess sample difficulty, then employs adaptive prompt refinement to provide varying levels of guidance. For problems the model can likely handle, GHPO primarily uses standard on-policy RL, encouraging exploration and self-discovery. But for more challenging samples, it seamlessly shifts to a form of imitation learning by offering explicit solution traces. This hybrid approach automatically balances exploration with direct guidance, preserving training efficiency for manageable tasks while effectively guiding the model through difficult ones, ultimately boosting both training stability and sample efficiency.
@@
  • We identify the critical role of capacity alignment and reward sparsity problem in RLVR and propose the GHPO framework to improve training stability and efficiency.
  • We introduce a novel framework called GHPO to detect sample difficulty and adaptively switch between on-policy reinforcement learning and guided imitation learning by adaptive prompt refinement.
@@ 3.1 From Static Guidance to a Dynamic Framework
As established, our core strategy is to integrate guidance directly into the reinforcement learning loop, conditioning the policy on partial ground-truth traces to overcome the reward sparsity detailed in Section 2.3. This approach is motivated by Assumption 1, which posits that such guidance increases the likelihood of success on difficult problems, thereby providing a valid learning signal where one would otherwise be absent. It’s worth noting that ground truth guidance, in the form of solution traces, is often available for most mathematics data. However, during the RLVR training process, this valuable solution trace information is typically overlooked in favor of only using the final ground truth answer.
@@ 4.3 Implementation Details
We conducted our experiments using the openr1 [17] codebase and the TRL [18] framework. GHPO training was performed on 8 powerful GPUs, each equipped with 80GB of memory and high memory bandwidth.